OpenStudy (anonymous):

√(2) + √(6) / √(8) + √(6) show work please

OpenStudy (mertsj):

simplify sqrt8 Then multiply numerator and denominator by the conjugate of the denominator.

OpenStudy (e.mccormick):

\[\frac{\sqrt{2}+\sqrt{6}}{\sqrt{8}+\sqrt{6}}\]or\[\sqrt{2}+\frac{\sqrt{6}}{\sqrt{8}}+\sqrt{6}\]

OpenStudy (zzr0ck3r):

\[\sqrt8 = 2\sqrt2\] so \[\frac{\sqrt2+\sqrt6}{2\sqrt2+\sqrt6}\]

OpenStudy (zzr0ck3r):

now do as @Mertsj said

OpenStudy (zzr0ck3r):

we want to get rid of the sqrts in the denom

OpenStudy (zzr0ck3r):

so \[\frac{\sqrt2+\sqrt6}{2\sqrt2+\sqrt6}*\frac{2\sqrt2-\sqrt6}{2\sqrt2-\sqrt6}\]